# Finely Stratified Rerandomization Designs

### v1

@@ 4.1 GMM Rerandomization
First, we generalize the imbalance metric $\mathcal{I}_{n}$ introduced in Definition 2.1, allowing rejection of a treatment allocation $D_{1:n}$ based on potentially nonlinear features of the in-sample distribution of treatments and covariates $(D_{i},X_{i})_{i=1}^{n}$. To define the nonlinear imbalance metric, let $m(X_{i},\beta)$ be a score function and define the within-arm GMM estimators $\widehat{\beta}_{1}$ and $\widehat{\beta}_{0}$ by
$\widehat{\beta}_{1}\in\operatorname*{argmin}_{\beta\in\mathbb{R}^{d_{\beta}}}|E_{n}[D_{i}m(X_{i},\beta)]|_{2}^{2},\quad\;\widehat{\beta}_{0}\in\operatorname*{argmin}_{\beta\in\mathbb{R}^{d_{\beta}}}|E_{n}[(1-D_{i})m(X_{i},\beta)]|_{2}^{2}.$ (4.1)
@@
Define $\mathcal{I}_{n}^{m}=\sqrt{n}(\widehat{\beta}_{1}-\widehat{\beta}_{0})$ as above, where $m(X,\beta)$ is a score satisfying Assumption 3.2. Suppose $d_{\beta}=d_{m}$ (exact identification) and let $A$ a symmetric acceptance region. Do the following: (1) form strata as in Definition 2.1. (2) Draw $D_{1:n}$ by stratified randomization. (3) If imbalance $\mathcal{I}_{n}^{m}=\sqrt{n}(\widehat{\beta}_{1}-\widehat{\beta}_{0})\in A$, accept. Otherwise, repeat from (2).
Observe that if $m(X_{i},\beta)=X_{i}-\beta$, then $\widehat{\beta}_{d}=\bar{X}_{d}$ and $\mathcal{I}_{n}^{m}=\mathcal{I}_{n}$, so linear rerandomization is a special case. Intuitively, the generalization $\mathcal{I}_{n}^{m}$ allows us to randomize until possibly nonlinear features of the covariates are balanced between the treatment and control groups.
@@
Let $f(X,\beta)$ be a parametric density model for the covariates $X$, which may be misspecified. Consider forming likelihood estimators $\widehat{\beta}_{1}\in\operatorname*{argmax}_{\beta}E_{n}[D_{i}\log f(X_{i},\beta)]$ and $\widehat{\beta}_{0}\in\operatorname*{argmax}_{\beta}E_{n}[(1-D_{i})\log f(X_{i},\beta)]$, then rerandomizing until the imbalance measure $\sqrt{n}|\widehat{\beta}_{1}-\widehat{\beta}_{0}|_{2}\leq\epsilon$. Under suitable regularity conditions, $\widehat{\beta}_{d}$ are GMM estimators as in Equation 4.1 with score function $m(X_{i},\beta)=\nabla_{\beta}f(X_{i},\beta)$, so this procedure is a GMM rerandomization with acceptance region $A=\{x:|x|_{2}\leq\epsilon\}$.
@@
Suppose $D_{1:n}$ is as in Definition 4.1 and Assumption 3.2 holds. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{\delta})+R$, independent RV’s with
$V_{\delta}=\min_{\eta\in\mathbb{R}^{d_{\beta}\times d_{\theta}}}v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})-\eta^{\prime}m(X_{i},\beta^{*})|\psi)].$
The residual $R\sim\eta_{0}^{\prime}Z_{m}\,|\,Z_{m}\in G_{m}A$ for $Z_{m}\sim\mathcal{N}(0,v_{D}^{-1}E[\operatorname{Var}(m(X_{i},\beta^{*})|\psi)])$.
Theorem 4.3 shows that by rerandomizing until $\sqrt{n}(\widehat{\beta}_{1}-\widehat{\beta}_{0})\in A$, we implicitly balance the influence function for the difference of GMM estimators in Equation 4.1, which depends both on $m(X_{i},\beta^{*})$ and the Jacobian matrix $G_{m}$.
This suggests an equivalent, but computationally simpler design with only one round of nonlinear estimation. In particular, let $\widehat{\beta}\in\operatorname*{argmin}_{\beta\in\mathbb{R}^{d_{\beta}}}|E_{n}[m(X_{i},\beta)]|_{2}^{2}$ and set rerandomization covariates $\widehat{h}_{i}=m(X_{i},\widehat{\beta})$, rerandomizing until $\sqrt{n}E_{n}[H_{i}\widehat{h}_{i}]\in G_{m}A$. This design is asymptotically equivalent to Definition 4.1, as shown in the next result.
@@
Suppose Assumption 3.1, 3.2 hold and let score $m(X,\beta)$ as in Definition 4.1. Let $D_{1:n}$ be rerandomized as in Definition 2.1 with $\widehat{h}_{i}=m(X_{i},\widehat{\beta})$ and acceptance region $G_{m}A$. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{\delta})+R$, with both variables identical to those in Theorem 4.3.
Corollary 4.4 shows that we can achieve the same effect as Definition 4.1 with a computationally simpler linear rerandomization that balances the estimated covariate $\widehat{h}_{i}=m(X_{i},\widehat{\beta})$. Next, we introduce a propensity score based approach that also attempts to balance nonlinear covariate features.
4.2 Propensity Rerandomization
To motivate a propensity based rerandomization procedure, note that under stratified randomization we have $E[D_{i}|W]=p$ for all units. In finite samples, however, the realized propensity $\widehat{p}(S)=E_{n}[D_{i}|X_{i}\in S]$ may be significantly different from $p$ in certain regions $S\subseteq\mathbb{R}^{d_{X}}$ of the covariate space. This implies that covariates are predictive of treatment assignments post-randomization, a form of “in-sample confounding.” To enforce balance, we could, for instance, reject allocations where $|\widehat{p}(S)-p|>\epsilon$ for some collection of sets $S$. To make this idea tractable without fully discretizing, consider a parametric propensity model $p(X,\beta)=L(X^{\prime}\beta)$ and define the MLE estimator
@@ 4.2 Propensity Rerandomization
We can measure the average gap between the estimated and true propensity score using the imbalance metric
@@
Intuitively, if $\mathcal{J}_{n}$ is large, then the covariates $X$ have power to predict treatment status in some parts of the covariate space. To avoid this, we propose rerandomizing until the imbalance metric $\mathcal{J}_{n}$ is below a threshold:
@@
Do the following: (1) form strata as in Definition 2.1. (2) Draw $D_{1:n}$ and estimate the propensity model in Equation 4.2. (3) If imbalance $\mathcal{J}_{n}\leq\epsilon$, accept. Otherwise, repeat from (2).
Before we state the main result, we require some regularity conditions
Assumption 4.6.
Impose the following conditions.
  1. (a)
    Let $L$ be twice differentiable, with $|L^{\prime}|_{\infty},|L^{\prime\prime}|_{\infty}<\infty$. For each $p\in(0,1)$, there is a unique $c$ with $L(c)=p$. Also, $|L^{\prime}(c)|>0$.
  2. (b)
    The score $m(D_{i},X_{i},\beta)=D_{i}\frac{L^{\prime}(X_{i}^{\prime}\beta)X_{i}}{L(X_{i}^{\prime}\beta)}-(1-D_{i})\frac{L^{\prime}(X_{i}^{\prime}\beta)X_{i}}{1-L(X_{i}^{\prime}\beta)}$ satisfies condition 3.2. The solution to Equation 4.2 exists and is unique.
  3. (c)
    Covariates $X=(1,h)$ for $E[|h|^{2}_{2}]<\infty$. Also, $E[\operatorname{Var}(h|\psi)]$, $E[XX^{\prime}]$ are full rank.
Condition (a) is satisfied by the logit and probit link functions. Our next result shows that propensity rerandomization as in Definition 4.5 is equivalent to a simpler linear rerandomization design, with an implicit choice of ellipsoidal acceptance region.
@@
Suppose $D_{1:n}$ is as in Definition 4.5. Require Assumptions 3.2, 4.6. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{\delta})+R$.
$V_{\delta}=\min_{\Gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}}v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})-\Gamma^{\prime}h|\psi)].$
@@
Theorem 4.7 shows that for any sufficiently regular link function, propensity rerandomization is asymptotically equivalent to the simpler quadratic rerandomization design in Example 2.2, with acceptance criterion $n(\bar{h}_{1}-\bar{h}_{0})^{\prime}\operatorname{Var}_{n}(h_{i})^{-1}(\bar{h}_{1}-\bar{h}_{0})\leq\epsilon$. Equivalently, propensity rerandomization behaves like linear rerandomization with $\mathcal{I}_{n}=\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})$ and ellipsoidal acceptance region $A=\operatorname{Var}(h)^{1/2}B(0,\epsilon)$. [^17] A related result was found by Ding and Zhao, (2024), who study rerandomizing until the p-value of a logistic regression coefficient is above a threshold.
Implicit Acceptance Regions. Both nonlinear designs in this section turned out to be equivalent to the standard rerandomization scheme in Definition 2.1, with a specific, implicit choice of rerandomization moments and acceptance region. However, note that the moments and acceptance region chosen by these procedures are entirely determined by the marginal covariate distribution. This implicit choice is not likely to be optimal, since the residual term $R_{A}\sim\Gamma_{0}^{\prime}Z_{h}|Z_{h}\in A$ depends not only on the covariates $Z_{h}\sim\mathcal{N}(0,v_{D}^{-1}E[\operatorname{Var}(h|\psi)])$, but also on the partially linear coefficient $\Gamma_{0}$. This coefficient is determined by the joint distribution of the assignment influence function $\Pi\delta(W,\theta_{0})$ and covariates $(\psi,h)$. In the next section, we show how to use prior information about this joint distribution to optimize the acceptance region and bound the variance of $R_{A}$.
@@ 5 Optimizing Acceptance Regions
In this section, we study efficient choice of the rerandomization acceptance region $A\subseteq\mathbb{R}^{d_{h}}$. For simplicity and intuition, we first restrict to the case of estimating $\theta_{n}=\operatorname{SATE}$, generalizing in what follows.
Imbalance Decomposition. The difference of means estimator $\widehat{\theta}=E_{n}[Y_{i}(1)-Y_{i}(0)]+E_{n}[H_{i}\bar{Y}_{i}]=\theta_{n}+E_{n}[H_{i}\bar{Y}_{i}]$ for $H_{i}=(D_{i}-p)/(p-p^{2})$. Intuitively, the scaled errors $\sqrt{n}(\widehat{\theta}-\theta_{n})=\sqrt{n}E_{n}[H_{i}\bar{Y}_{i}]$ are driven by imbalances in the outcome levels $\bar{Y}$ between treatment arms. The previous section showed that if covariates $h$ are predictive of $\bar{Y}$, we can reduce these imbalances by rerandomizing until $\sqrt{n}E_{n}[H_{i}h_{i}]=\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})\in A$. To study the role of the acceptance region $A$, let $(\gamma_{0},t_{0})$ be solutions to the partially linear prediction problem in Equation 3.3 and consider the expansion [^18] This is without loss of generality. Note that we do not impose well-specification $E[e|\psi,h]=0$.
@@
We use this to decompose the imbalance in outcome levels $\bar{Y}$ into imbalances in the covariates $\psi$ and $h$ and residuals $e$. In particular, we can now write the imbalance decomposition
$E_{n}[H_{i}\bar{Y}_{i}]=E_{n}[H_{i}t_{0}(\psi_{i})]+\gamma_{0}^{\prime}E_{n}[H_{i}h_{i}]+E_{n}[H_{i}e_{i}]\equiv I_{1}+I_{2}+I_{3}.$
The analysis in Section 3 showed that for any acceptance region $A\subseteq\mathbb{R}^{d_{h}}$:
  1. (1)
[…]
  2. (2)
    The components $I_{2}+I_{3}=\gamma_{0}^{\prime}\sqrt{n}E_{n}[H_{i}h_{i}]+\sqrt{n}E_{n}[H_{i}e_{i}]\Rightarrow R_{A}+v_{D}^{-1/2}\mathcal{N}(0,\operatorname{Var}(e))$ are asymptotically independent, with $\operatorname{Var}(e)$ not depending on $A$.
In particular, it suffices to choose $A$ to minimize the component $I_{2}=\gamma_{0}^{\prime}\sqrt{n}E_{n}[H_{i}h_{i}]$. This suggests an oracle acceptance criterion, rerandomizing until $|\gamma_{0}^{\prime}\sqrt{n}E_{n}[H_{i}h_{i}]|\leq\epsilon$, with acceptance region $A=\{a:|\gamma_{0}^{\prime}a|\leq\epsilon\}$. However, this acceptance region is infeasible since $\gamma_{0}$ is unknown at design-time. Instead, we take a minimax approach, allowing the researcher to incorporate prior information about $\gamma_{0}$.

### v2

@@ 4.1 GMM Rerandomization
First, we generalize the imbalance metric $\mathcal{I}_{n}$ introduced in Definition 2.1, allowing rejection of a treatment allocation $D_{1:n}$ based on potentially nonlinear features of the in-sample distribution of treatments and covariates $(D_{i},X_{i})_{i=1}^{n}$. We can define a large class of nonlinear imbalance metrics by letting $m(X_{i},\beta)$ be a score function and considering within-arm GMM estimators $\widehat{\beta}_{1}$ and $\widehat{\beta}_{0}$ defined by
$E_{n}[D_{i}m(X_{i},\widehat{\beta}_{1})]=0,\quad\;E_{n}[(1-D_{i})m(X_{i},\widehat{\beta}_{0})]=0.$ (4.1)
We propose to rerandomize until the within-arm parameter estimates are approximately equal, $\sqrt{n}(\widehat{\beta}_{1}-\widehat{\beta}_{0})\approx 0$.
@@
Define $\mathcal{I}_{n}^{m}=\sqrt{n}(\widehat{\beta}_{1}-\widehat{\beta}_{0})$ as above, where $m(X,\beta)$ is a score satisfying Assumption 3.2. Suppose $d_{\beta}=d_{m}$ (exact identification) and let $A$ be a symmetric acceptance region. Do the following: (1) form groups as in Definition 2.1. (2) Draw $D_{1:n}$ by stratified randomization. (3) If imbalance $\mathcal{I}_{n}^{m}=\sqrt{n}(\widehat{\beta}_{1}-\widehat{\beta}_{0})\in A$, accept $D_{1:n}$. Otherwise, repeat from (2).
Intuitively, the generalized imbalance metric $\mathcal{I}_{n}^{m}$ allows us to randomize until possibly nonlinear features of the covariates are balanced between the treatment and control groups. Observe that if $m(X_{i},\beta)=X_{i}-\beta$, then $\widehat{\beta}_{d}=\bar{X}_{d}$ for $d=0,1$ and $\mathcal{I}_{n}^{m}=\mathcal{I}_{n}$, so linear rerandomization is a special case.
@@
Let $f(X,\beta)$ be a parametric density model for covariates $X$, which may be misspecified. After drawing $D_{1:n}$ by stratified randomization, consider forming (quasi) maximum likelihood estimators $\widehat{\beta}_{1}\in\operatorname*{argmax}_{\beta}E_{n}[D_{i}\log f(X_{i},\beta)]$ and $\widehat{\beta}_{0}\in\operatorname*{argmax}_{\beta}E_{n}[(1-D_{i})\log f(X_{i},\beta)]$ for the density of covariates assigned to each treatment arm, rerandomizing until the estimated parameters $\sqrt{n}|\widehat{\beta}_{1}-\widehat{\beta}_{0}|_{2}\leq\epsilon$. Under regularity conditions, [^17] For example, if $\beta\to\log f(X,\beta)$ is a.s. strictly concave, the key identification condition in Assumption 3.2 will be satisfied. $\widehat{\beta}_{d}$ are GMM estimators as in Equation 4.1 with score function $m(X_{i},\beta)=\nabla_{\beta}\log f(X_{i},\beta)$, so this procedure is a GMM rerandomization with acceptance region $A=\{x:|x|_{2}\leq\epsilon\}$.
@@
Suppose $D_{1:n}$ is as in Definition 4.1 and Assumption 3.2 holds. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})+R$, independent RV’s with
$V_{a}=\min_{\gamma\in\mathbb{R}^{d_{m}\times d_{\theta}}}v_{D}^{-1}E[\operatorname{Var}(\Pi a(W,\theta_{0})-\gamma^{\prime}m(X_{i},\beta^{*})|\psi)].$ (4.2)
The residual $R\sim\gamma_{0}^{\prime}Z_{m}\,|\,Z_{m}\in G_{m}A$ for $Z_{m}\sim\mathcal{N}(0,v_{D}^{-1}E[\operatorname{Var}(m(X_{i},\beta^{*})|\psi)])$, where $\gamma_{0}$ is optimal in Equation 4.2.
Theorem 4.3 shows that by rerandomizing until $\sqrt{n}(\widehat{\beta}_{1}-\widehat{\beta}_{0})\in A$, we implicitly balance the influence function $-G_{m}^{-1}m(X_{i},\beta^{*})$ for the difference of GMM estimators in Equation 4.1. This suggests an equivalent, but computationally much simpler design with only one round of nonlinear estimation. In particular, let $\widehat{\beta}$ solve $E_{n}[m(X_{i},\widehat{\beta})]=0$ be the pooled GMM estimator and set rerandomization covariates $\widehat{h}_{i}=m(X_{i},\widehat{\beta})$, rerandomizing until $\sqrt{n}E_{n}[H_{i}\widehat{h}_{i}]\in G_{m}A$. The next result shows that this design, which generalizes Definition 2.1 to allow for estimated covariates, is asymptotically equivalent to the GMM rerandomization in Definition 4.1.
@@
Suppose Assumption 3.1, 3.2 hold and let $m(X,\beta)$ be as in Definition 4.1. Let $D_{1:n}$ be rerandomized as in Definition 2.1 with $\widehat{h}_{i}=m(X_{i},\widehat{\beta})$ and acceptance region $G_{m}A$. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})+R$, with both variables identical to those in Theorem 4.3.
Corollary 4.4 is a useful tool for showing the equivalence of computationally intensive designs based on nonlinear estimation with simpler linear schemes. For instance, Example 4.5 shows that density rerandomization over an exponential family with sufficient statistic $r(X_{i})$ is asymptotically equivalent to linear rerandomization with $h_{i}=r(X_{i})$.
Example 4.5 (Density Rerandomization).
[…]
$f(x,\beta)=\exp\left(\sum_{j}x_{j}\beta_{j}+\sum_{j<l}x_{j}x_{l}\beta_{jl}-t(\beta)\right).$
The sufficient statistic is $r(x)=((x_{j})_{j},(x_{j}x_{l})_{j<l})$, and the parameters $\beta_{jl}$ model correlation between the binary variables $x_{j}$ and $x_{l}$. By the discussion above, a design that rerandomizes based on the difference of quasi-MLE density estimates in this family [^20] This is well-motivated when $\psi$ is expected to be more important than $(x_{j})_{j}$. We don’t want to stratify on both, since this could radically decrease match quality on $\psi$. is asymptotically equivalent to the much simpler linear rerandomization in Definition 2.1 with covariates $h_{i}=((x_{j})_{j},(x_{j}x_{l})_{j<l})$.
4.2 Propensity Score Rerandomization
To motivate a propensity score based rerandomization procedure, note that under stratified randomization we have $E[D_{i}|X_{i}]=p$ for all units. In finite samples, however, the realized propensity $\widehat{p}(B)=E_{n}[D_{i}|X_{i}\in B]$ may significantly diverge from $p$ in certain regions $B\subseteq\mathbb{R}^{d_{X}}$ of the covariate space. This implies that covariates are predictive of treatment assignments post-randomization, a form of “in-sample confounding,” which of vanishes as $n\to\infty$ but affects precision. To prevent this, we could, for instance, reject allocations where $|\widehat{p}(B)-p|>\epsilon$ for some collection of sets $B$. To make this idea tractable without fully discretizing, consider a parametric propensity model $p(X,\beta)=L(X^{\prime}\beta)$ and define the MLE estimator
@@ 4.2 Propensity Score Rerandomization
We can measure the average gap between the estimated and true propensity score using
@@
Intuitively, if $\mathcal{J}_{n}$ is large, then the covariates $X$ are predictive of treatment status in some parts of the covariate space. To avoid this, we propose rerandomizing until the imbalance metric $\mathcal{J}_{n}$ is below a threshold:
@@
Do the following: (1) form groups as in Definition 2.1. (2) Draw $D_{1:n}$ and estimate the propensity model in Equation 4.3. (3) If imbalance $\mathcal{J}_{n}\leq\epsilon$, accept. Otherwise, repeat from (2).
Refer to caption Refer to caption
Figure 1: Propensity rerandomization (Definition 4.6) with $p=1/2$ for $Z\sim\operatorname{Unif}[0,1]$ and $X=B(Z)$ a B-spline basis. LHS: $D_{1:n}$ and estimated propensity with $\widehat{p}(Z)\ll 1/2$, for $Z\in[0.4,0.9]$. RHS: Accepted allocation $D_{1:n}$ with $\mathcal{J}_{n}\leq\epsilon$
Our next result shows that propensity rerandomization as in Definition 4.6 is equivalent to a simpler linear rerandomization design, with an implicit choice of ellipsoidal acceptance region. We require some extra regularity conditions on the link function $L$, which for brevity we state in Appendix 8.5.
@@
Suppose $D_{1:n}$ is as in Definition 4.6. Require Assumptions 3.2, 8.12. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})+R$.
$V_{a}=\min_{\gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}}v_{D}^{-1}E[\operatorname{Var}(\Pi a(W,\theta_{0})-\gamma^{\prime}h|\psi)].$
@@
Theorem 4.7 shows that for any sufficiently regular link function, propensity rerandomization is asymptotically equivalent to the quadratic rerandomization design in Example 2.2, with acceptance criterion $n(\bar{h}_{1}-\bar{h}_{0})^{\prime}\operatorname{Var}_{n}(h_{i})^{-1}(\bar{h}_{1}-\bar{h}_{0})\leq\epsilon$. Equivalently, propensity rerandomization behaves like linear rerandomization with $\mathcal{I}_{n}=\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})$ and ellipsoidal acceptance region $A=\operatorname{Var}(h)^{1/2}B(0,\epsilon)$. [^21] A related result was found by Ding and Zhao, (2024), who study rerandomizing until the p-value of a logistic regression coefficient is above a threshold.
Implicit Acceptance Regions. Both nonlinear designs in this section turned out to be equivalent to the standard rerandomization scheme in Definition 2.1, with a specific, implicit choice of rerandomization moments and acceptance region determined by the choice of score $m$ and marginal covariate distribution. However, this implicit choice is not likely to be optimal, since the residual term in the asymptotic error distribution $R_{A}\sim\gamma_{0}^{\prime}Z_{h}|Z_{h}\in A$ depends on both the covariates $Z_{h}\sim\mathcal{N}(0,v_{D}^{-1}E[\operatorname{Var}(h|\psi)])$, and the partially linear coefficient $\gamma_{0}$. This coefficient is determined by the joint distribution of the assignment influence function $\Pi a(W,\theta_{0})$ and covariates $(\psi,h)$. In the next section, we show how to use prior information about this joint distribution to optimize the acceptance region and bound the variance of $R_{A}$.
@@ 5 Optimizing Acceptance Regions
In this section, we study efficient choice of the acceptance region $A\subseteq\mathbb{R}^{d_{h}}$. We propose a novel minimax rerandomization criterion that minimizes asymptotic computational cost subject to a strict lower bound on statistical efficiency.
For intuition, we first restrict to the case of estimating $\theta_{n}=\operatorname{SATE}$, generalizing in what follows. Example 3.9 showed that $\sqrt{n}(\widehat{\theta}-\operatorname{SATE})|W_{1:n}\Rightarrow\mathcal{N}(0,V(\gamma_{0}))+\gamma_{0}^{\prime}Z_{hA}$, independent RV’s with $Z_{hA}=Z_{h}|Z_{h}\in A$ and $V(\gamma_{0})$ that does not depend on $A$. The coefficient $\gamma_{0}$ is identified by the partially linear regression [^22] This expansion is without loss of generality. We do not impose well-specification $E[e|\psi,h]=0$.
@@
The term $R_{A}=\gamma_{0}^{\prime}Z_{hA}$ arises from fluctuations of $\bar{Y}$ predictable by $h$ that are not balanced by rerandomization due to slackness in the acceptance region, $A\not=\{0\}$. More precisely, $\gamma_{0}^{\prime}Z_{hA}$ is the rerandomized limiting distribution of $\gamma_{0}^{\prime}\mathcal{I}_{n}=\gamma_{0}^{\prime}\sqrt{n}(\bar{h}_{1}-\bar{h}_{0})$, the projection of covariate imbalances in $h$ along the direction $\gamma_{0}$. This suggests an oracle acceptance criterion that rerandomizes until $|\gamma_{0}^{\prime}\mathcal{I}_{n}|\leq\epsilon$, with acceptance region $A=\{x:|\gamma_{0}^{\prime}x|\leq\epsilon\}$. Of course, this is infeasible since $\gamma_{0}$ is unknown when designing the experiment.